# MEMoE: Enhancing Model Editing with Mixture of Experts Adaptors

### v1

@@ Appendix A Limitation
Secondly, the excellent performance of MOMoE indicates exciting potential for practical applications of model editing technology in specific domain such as medicine and education. However, this study is limited to testing on mainstream model editing datasets. Future work could involve performance testing on domain-specific datasets to further advance the application of model editing technology. Additionally, model editing techniques could be applied to various types of tasks. Specifically, beyond factual knowledge editing, they can be used to eliminate hallucinations, biases, and privacy information. However, our experiments focus only on general editing tasks (such as batch editing and sequential batch editing), which are relatively well-studied and universally evaluated in model editing, and do not address issues like mitigating hallucinations.
@@ Appendix D Baselines and Implementation Details
MOMoE
For main experiments in §4.2 and §4.3, the experts number is 4 and $top_{k}$ value is 1 for MOMoE, which has yielded the best experimental results within the current computational resources. And the modification is applied to layer 16 for LLaMA2-7B and layer 18 for GPT2-XL (consistent with the findings of ROME [33]). We also adopt auxiliary loss for balancing the top-k selection of routing following Switch Transformers [11]. We employ the NLTK [^1] https://www.nltk.org tool to extract named entities from input text. Furthermore, in cases where there are multiple named entities present in the input, we utilize the average pooling of the embeddings of these entities as the knowledge anchor [49]. The experiment is deployed on NVIDIA RTX 3090 Tensor Core GPUs, and we use 4 GPUs on training and single GPU on evaluation. The training time for the batch editing task is approximately two minutes, while the training time for the sequential editing task is about one hour (learning rate for all experiments is $2e^{-4}$). The implementation details of baselines are as follow [29]:
@@ E.2 Different Model Configurations
We provide results under different expert number, target layer and top-k settings in Figure LABEL:table:additional_results. The experimental results demonstrate that under various parameter settings, MOMoE consistently exhibits outstanding overall performance, thoroughly illustrating the superiority of its method. Furthermore, the minimal fluctuation in the scores of the model’s reliability and locality score confirms that the accuracy and locality performance are inherently guaranteed by the characteristics of the bypass MoE structure, which is consistent with the design principles of MOMoE discussed in §3.1.
Table 6: Additional experimental results of different MOMoE configurations.

### v2

@@ Appendix A Limitation
Secondly, the excellent performance of MEMoE indicates exciting potential for practical applications of model editing technology in specific domain such as medicine and education. However, this study is limited to testing on mainstream model editing datasets. Future work could involve performance testing on domain-specific datasets to further advance the application of model editing technology. Additionally, model editing techniques could be applied to various types of tasks. Specifically, beyond factual knowledge editing, they can be used to eliminate hallucinations, biases, and privacy information. However, our experiments focus only on general editing tasks (such as batch editing and sequential batch editing), which are relatively well-studied and universally evaluated in model editing, and do not address issues like mitigating hallucinations.
@@ Appendix D Baselines and Implementation Details
MEMoE
For main experiments in §4.2 and §4.3, the experts number is 4 and $top_{k}$ value is 1 for MEMoE, which has yielded the best experimental results within the current computational resources. And the modification is applied to layer 16 for LLaMA2-7B and layer 18 for GPT2-XL (consistent with the findings of ROME [33]). We also adopt auxiliary loss for balancing the top-k selection of routing following Switch Transformers [11]. We employ the NLTK [^1] https://www.nltk.org tool to extract named entities from input text. Furthermore, in cases where there are multiple named entities present in the input, we utilize the average pooling of the embeddings of these entities as the knowledge anchor [49]. The experiment is deployed on NVIDIA RTX 3090 Tensor Core GPUs, and we use 4 GPUs on training and single GPU on evaluation. The training time for the batch editing task is approximately two minutes, while the training time for the sequential editing task is about one hour (learning rate for all experiments is $2e^{-4}$). The implementation details of baselines are as follow [29]:
@@ E.2 Different Model Configurations
We provide results under different expert number, target layer and top-k settings in Figure LABEL:table:additional_results. The experimental results demonstrate that under various parameter settings, MEMoE consistently exhibits outstanding overall performance, thoroughly illustrating the superiority of its method. Furthermore, the minimal fluctuation in the scores of the model’s reliability and locality score confirms that the accuracy and locality performance are inherently guaranteed by the characteristics of the bypass MoE structure, which is consistent with the design principles of MEMoE discussed in §3.1.
Table 6: Additional experimental results of different MEMoE configurations.